# Off-Road Navigation via Implicit Neural Representation of Terrain Traversability

### v2

@@ II-B1 Path Planner
We use A [hart1968formal] to generate an initial path. To construct the A cost map, we first query $f_{\texttt{g\_max}}$ at grid locations sampled at a specified coarse resolution. This efficiently produces a coarse geometric grid, from which we compute two metrics: slope (the spatial gradient of $f_{\texttt{g\_max}}$) and step size (the elevation difference between adjacent cells). These metrics are normalized to $[0,1]$ using the maximum acceptable slope and step size respectively. We take the product of the two metrics to fuse them into a single cost map. In our A implementation, the traversal cost of an edge between neighboring cells is given by the average cell cost multiplied by the Euclidean distance between cell centers. The heuristic is chosen as the straight-line distance to the goal multiplied by the minimum cell cost in the map. Because any feasible path must be at least as long as the Euclidean distance and cannot have a per-meter cost lower than this minimum, the heuristic never overestimates the true cost-to-go and is therefore admissible, while still providing a strong guidance signal for the search.
Figure 4: Inference time for the explicit decoder and implicit decoder vs. number of query points. Note that the explicit decoder outputs a fixed-size $100\times 100$ grid map. Therefore, $10^{4}$ query points to the implicit decoder is equivalent to reconstructing a grid map with the same grid resolution as the explicit decoder. However, for planning and trajectory optimization, the required number of queries is usually far smaller than the full grid size.
II-B2 Trajectory Optimization
@@
Given an initial path $\{p^{\text{init}}_{i}=(p^{\text{init}}_{i,x},p^{\text{init}}_{i,y})\}_{i=1}^{N_{p}}$ in the local map frame, obtained from A, we refine it via gradient–based trajectory optimization to minimize travel time and vertical acceleration while ensuring smoothness and dynamic feasibility. We parameterize the geometric path by a set of Catmull–Rom [catmull1974class] control points $\mathcal{C}=\{c_{m}\}_{m=1}^{M}$ with fixed endpoints $c_{1}=p^{\text{init}}_{1}$, $c_{M}=p^{\text{init}}_{N_{p}}$ and trainable midpoints $\{c_{2},\dots,c_{M-1}\}$. Centripetal Catmull–Rom interpolation [yuksel2009parameterization] yields a dense path $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ ($N>N_{p}$) sampled along arc length.
@@
We also define a curvature-limited speed cap for each segment $i$ as the smooth minimum [nesterov2005smooth] of two speed caps:
$v_{\text{cap},i}$ $=\operatorname{smin}_{\tau}\!\Bigl(v_{\max},\,\sqrt{\tfrac{a_{\text{lat}}^{\text{max}}}{|\kappa_{i}|+\varepsilon}}\Bigr)$
$=-\tau\log(e^{-v_{\max}/\tau}+e^{-\left(\sqrt{\tfrac{a_{\text{lat}}^{\text{max}}}{|\kappa_{i}|+\varepsilon}}\right)/\tau}),$
@@
$v_{\text{pref},i}=\operatorname*{arg\,min}_{v>0}\frac{w_{\text{time}}}{v}+w_{\text{bump}}b^{\alpha}_{i}v,$
where $w_{\text{time}},w_{\text{bump}}$ are weights used to balance trade-offs between travel time and predicted vertical acceleration. The first term in the objective encourages shorter travel time, while the second term penalizes high vertical acceleration, taking into account that vertical acceleration scales proportionally to vehicle speed multiplied by the bumpiness estimate. The exponent $\alpha\in\mathbb{N}$ in $b^{\alpha}_{i}$ helps rescales the bumpiness prediction. Empirically, setting $\alpha$ to a higher value (e.g. $2$ or $3$ vs. $1$) creates a larger contrast in the speed profile (i.e. slowing down more in bumpy regions and speeding up more in smooth regions). Note that the objective is convex in $v$ and the minimizer can be obtained by setting its gradient to 0, which yields $v_{\text{pref},i}=\sqrt{\frac{w_{\text{time}}}{w_{\text{bump}}\,(b^{\alpha}_{i}+\varepsilon)}}.$ As before, to ensure differentiability, we fuse the speed cap and the preferred speed using the smooth minimum operator $v_{i}=\operatorname{smin}_{\tau}(v_{\text{cap},i},\,v_{\text{pref},i})$. Let $\bar{v}_{i}=\tfrac{1}{2}(v_{i}+v_{i+1})$ denote the segment speed and $\bar{\kappa}_{i}=\tfrac{1}{2}(\kappa_{i}+\kappa_{i+1})$ denote the segment curvature. Our objective combines a time surrogate, a speed-weighted bump cost, and geometric regularizers:
@@
where $\lambda_{b},\lambda_{s},\lambda_{\kappa}$ are weights used to balance different terms. Note that the objective $\mathcal{J}(\mathcal{C})$ is differentiable with respect to $\mathbf{P}$ thanks to the differentiability of $f_{\texttt{bump}}$ with respect to the query points as well as the differentiability of $\bar{v}_{i},\bar{\kappa}_{i},\Delta s_{i}$ with respect to $\mathbf{P}$. Since the interpolation from $\mathcal{C}$ to $\mathbf{P}$ is differentiable, the objective is differentiable with respect to $\mathcal{C}$, thus enabling gradient-based optimization. We optimize only the interior control points using $\operatorname{Adam}$ [kingma2014adam] with gradient clipping, and project them onto the local planning bounds $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$. Start and goal (i.e. $c_{1},c_{M}$) are held fixed.
After the optimization, we perform a 1-D time-scaling along arc length to obtain a dynamically feasible velocity profile following standard trajectory time-parameterization approaches [bobrow1985time, pham2018new]. We enforce $v\leq v_{\max}$, $v^{2}|\kappa|\leq a_{\text{lat}}^{\text{max}}$, and tangential acceleration limits $a\in[-a_{\text{dec}},\,a_{\text{acc}}]$, while retaining the same soft trade-off between time and bump cost to shape the speed profile. Finally, we recover yaw from the path tangent and angular rate via $\omega=v\,\kappa$, yielding the outputs $\{\,t,\,\mathbf{P},\,\mathrm{yaw}(t),\,v(t),\,\omega(t)\}$.
II-B3 Tracking Controller
We formulate an MPC problem to track the optimized trajectory. Let $x=[p_{x},p_{y},\theta]^{T},u=[v,\omega]^{T}$, where $p_{x},p_{y}$ denote the $x,y$ coordinates of the vehicle, $\theta$ denotes the yaw angle, and $v,\omega$ demote the linear speed and angular rate. The optimization problem is then:
@@ II-B3 Tracking Controller
which is solved using do-mpc [fiedler2023mpc, andersson2019casadi, wachter2006implementation].
III Simulation Experiments
@@
III-A Baselines
We compare against one geometry-based and one traction-based baseline. For the geometry-based method, we implement a RoadRunner-style method [frey2024roadrunner]: geometric risk factors (step size, slope, curvature) are fused via Conditional Value at Risk (CVaR) to form a risk map, which is then used by an MPPI controller. Our implementation uses our encoder and the WayFASTER decoder [gasparino2024wayfaster] due to lack of publicly available code. We refer to this baseline as Geometry. For the traction-based baseline, we use WayFASTER [gasparino2024wayfaster] with its released network architecture [^1] https://github.com/matval/wayfaster, paired with an MPPI controller following the formulation in the original paper. We implemented both MPPI controllers by adapting a public PyTorch-based codebase [^2] https://github.com/UM-ARM-Lab/pytorch_mppi. We use a horzion $N=20$ with a time step of $dt=0.1$\mathrm{s}$\$.
Figure 5: Grassland scenario. The robot’s forward direction ($x$) and leftward direction ($y$) are represented by the red and green axes, respectively. Examples of paths produced by Geo-20 (blue), WF-20 (orange), and TRAIL (green) are shown on the left. The bottom left corner shows the camera image from the robot. On the right side, the top shows the initial path generated by A overlayed on the geometric risk map. Bottom right shows the optimized trajectory overlayed on the predicted bumpiness value. The trajectory is color-coded by speed.
III-B Dataset and Training
@@ III-C Metric
III-D Timing Analysis
Timings are performed on a workstation with an RTX 4090 GPU, an AMD Ryzen 9 7950x 16-core processor with 32 threads, and 128 GB of RAM. A detailed analysis of decoder timing is shown in Figure 4. The encoder of our model and Geometry take $25$\mathrm{ms}$\$ per inference and WayFASTER’s encoder takes $20$\mathrm{ms}$\$. With horizon $N=20$, Geometry’s MPPI takes $10$\mathrm{ms}$\$ per inference and WayFASTER’s MPPI takes $25$\mathrm{ms}$\$ per inference. For the proposed method, A takes $5~$\mathrm{ms}$\$ (to generate a path with 40 points), trajectory optimization takes $7.5~$\mathrm{ms}$\$ per iteration (to optimize an initial path with 30 points, which are downsampled from the A path), and the MPC take $6$\mathrm{ms}$\$ with horizon $N=20$.
Figure 6: Forest scenario. Each row shows one set of experiments. The first column shows Path 1 and the second column shows Path 2 of each set of experiments. The forest consists of tall grass, bushes, rocks, logs, and trees. Images are tone- and color-corrected for visibility. The proposed method is able to navigate efficiently while the two baselines struggle: Geometry is conservative due to the density of objects with high elevations while WayFASTER is too optimistic since there are no collision events with similar objects in the training data. On the other hand, our method is able to recognize and avoid rocks while not being constrained by the tall grass.
III-E Navigation Performance
@@
Grassland The Grassland scenario (see Figure 5) features a rock hidden in tall grass (bottom left of Figure 5). The quantitative results in Table I show that Geometry, whose inflated geometric map cannot distinguish the rock from tall grass, produces an almost uniform risk map, drives straight toward the goal, runs over the rock, and incurs large vertical acceleration (Geo-20). Because CVaR values lie in $[0,1]$ [frey2024roadrunner], MPPI’s behavior is highly sensitive to the trade-off between risk and goal distance: high risk weight leads to oscillations and failure to reach the goal, while low risk weight reproduces Geometry’s straight-line behavior, consistent with our ablation study in Section III-F. WayFASTER detects the rock but is overly optimistic, briefly clipping it and producing high vertical acceleration (WF-20), likely due to weak BEV cues from sparse supervision and a limited encoder. In contrast, our method maintains clearance, using bumpiness during trajectory optimization to push the path away from the rock while remaining less conservative in grass, yielding a smoother, shorter trajectory and an adaptive speed profile that slows near the rock and accelerates through grass. Increasing the MPPI horizon from $20$ to $50$ confirms that the baselines are not horizon-limited: Geo-50 becomes indecisive and slower due to accumulated risk, while WF-50 matches WF-20 (the single WF-20 failure resulted from odometry divergence). Thus, we use $N=20$ for subsequent experiments.
@@ III-F Ablation Study
To evaluate the proposed planner, we conduct an ablation study by generating fixed-resolution cost maps from our trained model and using MPPI as the planner, following common practice. We implement five variants:
@@
In summary, each MPPI formulation minimizes the distance to goal and cost along the path, with a small penalty on control effort. MPPI-A-Bump has an additional cost term that penalizes deviations from the initial path. The planning horizon is set to $N=20$ with $dt=0.1s$ for all variants.
@@
IV Field Experiments
We further compare TRAIL with the two baseline methods in real-world hardware experiments. We used an AgileX Scout Mini equipped with a ZED 2i stereo camera and an Ouster OS1-32-GEN2.0 LiDAR. The training data consists of less than 2000 samples, which were all obtained around an area similar to the testing area but with a Clearpath Husky A300 UGV that has the same sensor stack as the Scout. The training data consists of the Husky driving over dirt, paved road, tall grass, tree roots, and branches on the ground, but with no purposeful collision with trees. During data collection, we observed significant odometry drift in the data collected using the Scout Mini (despite post-processing and parameter tuning) due to noisy sensor readings caused by the challenging terrain, which is not ideal for generating training labels. In contrast, the Husky has smoother sensor readings due to its greater mass and better traction, which helps reduces odometry drift. Note that this introduces a distribution shift during testing, We used KISS ICP [vizzo2023kiss] to generate the odometry used in training data processing, since we found that it generated better odometry during offline processing with minimal tuning, and DLIO [chen2022direct] to provide odometry at a higher rate during field deployment.

### v3

@@ III-B1 Path Planner
We use A [19] to generate an initial path. To construct the A cost map, we first query $f_{\texttt{g\_max}}$ at grid locations sampled at a specified coarse resolution. This efficiently produces a coarse geometric grid, from which we compute two metrics: slope (the spatial gradient of $f_{\texttt{g\_max}}$) and step size (the elevation difference between adjacent cells). These metrics are normalized to $[0,1]$ using the maximum acceptable slope and step size respectively. We take the product of the two metrics to fuse them into a single cost map. In our A implementation, the traversal cost of an edge between neighboring cells is given by the average cell cost multiplied by the Euclidean distance between cell centers. The heuristic is chosen as the straight-line distance to the goal multiplied by the minimum cell cost in the map. Because any feasible path must be at least as long as the Euclidean distance and cannot have a per-meter cost lower than this minimum, the heuristic never overestimates the true cost-to-go and is therefore admissible, while still providing a strong guidance signal for the search.
III-B2 Trajectory Optimization
@@
Given an initial path $\{p^{\text{init}}_{i}=(p^{\text{init}}_{i,x},p^{\text{init}}_{i,y})\}_{i=1}^{N_{p}}$ in the local map frame, obtained from A, we refine it via gradient–based trajectory optimization to minimize travel time and vertical acceleration while ensuring smoothness and dynamic feasibility. We parameterize the geometric path by a set of Catmull–Rom [8] control points $\mathcal{C}=\{c_{m}\}_{m=1}^{M}$ with fixed endpoints $c_{1}=p^{\text{init}}_{1}$, $c_{M}=p^{\text{init}}_{N_{p}}$ and trainable midpoints $\{c_{2},\dots,c_{M-1}\}$. Centripetal Catmull–Rom interpolation [52] yields a dense path $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ ($N>N_{p}$) sampled along arc length.
@@
We also define a curvature-limited speed cap for each segment $i$ as the smooth minimum [33] of two speed caps:
$v_{\text{cap},i}$ $=\operatorname{smin}_{\tau}\!\Bigl(v_{\max},\,\sqrt{\tfrac{a_{\text{lat}}^{\text{max}}}{|\kappa_{i}|+\varepsilon}}\Bigr)$ (3)
$=-\tau\log\!\left(e^{-v_{\max}/\tau}+e^{-\left(\sqrt{\tfrac{a_{\text{lat}}^{\text{max}}}{|\kappa_{i}|+\varepsilon}}\right)/\tau}\right).$
@@
$v_{\text{pref},i}=\arg\min_{v>0}\frac{w_{\text{time}}}{v}+w_{\text{bump}}\,b^{\alpha}_{i}\,v$ (4)
where $w_{\text{time}},w_{\text{bump}}$ are weights used to balance trade-offs between travel time and predicted vertical acceleration. The first term in the objective encourages shorter travel time, while the second term penalizes high vertical acceleration, taking into account that vertical acceleration scales proportionally to vehicle speed multiplied by the bumpiness estimate. The exponent $\alpha\in\mathbb{N}$ in $b^{\alpha}_{i}$ helps rescale the bumpiness prediction. Empirically, setting $\alpha$ to a high value (e.g. $3$ vs. $1$) creates a larger contrast in the speed profile (i.e. slowing down more in bumpy regions and speeding up more in smooth regions). Note that the objective is convex in $v$ and the minimizer can be obtained by setting its gradient to 0, which yields $v_{\text{pref},i}=\sqrt{\frac{w_{\text{time}}}{w_{\text{bump}}\,(b^{\alpha}_{i}+\varepsilon)}}.$ As before, to ensure differentiability, we fuse the speed cap and the preferred speed using the smooth minimum operator $v_{i}=\operatorname{smin}_{\tau}(v_{\text{cap},i},\,v_{\text{pref},i})$. Let $\bar{v}_{i}=\tfrac{1}{2}(v_{i}+v_{i+1})$ denote the segment speed and $\bar{\kappa}_{i}=\tfrac{1}{2}(\kappa_{i}+\kappa_{i+1})$ denote the segment curvature. Our objective combines a time surrogate, a speed-weighted bump cost, and geometric regularizers:
@@
where the weights $\lambda_{b},\lambda_{s},\lambda_{\kappa}$ balance the terms. Note that the objective $\mathcal{J}(\mathcal{C})$ is differentiable with respect to $\mathbf{P}$ thanks to the differentiability of $f_{\texttt{bump}}$ with respect to the query points as well as the differentiability of $\bar{v}_{i},\bar{\kappa}_{i},\Delta s_{i}$ with respect to $\mathbf{P}$. Since the interpolation from $\mathcal{C}$ to $\mathbf{P}$ is differentiable, the objective is differentiable with respect to $\mathcal{C}$, thus enabling gradient-based optimization. We optimize only the interior control points using $\operatorname{Adam}$ [22] with gradient clipping, and project them onto the local planning bounds $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$. Start and goal (i.e. $c_{1},c_{M}$) are held fixed.
After the optimization, we perform a 1-D time-scaling along arc length to obtain a dynamically feasible velocity profile following standard trajectory time-parameterization approaches [2, 38]. We enforce $v\leq v_{\max}$, $v^{2}|\kappa|\leq a_{\text{lat}}^{\text{max}}$, and tangential acceleration limits $a\in[-a_{\text{dec}},\,a_{\text{acc}}]$, while retaining the same soft trade-off between time and bump cost to shape the speed profile. Finally, we recover yaw from the path tangent and angular rate via $\omega=v\,\kappa$, yielding the outputs $\{\,t,\,\mathbf{P},\,\mathrm{yaw}(t),\,v(t),\,\omega(t)\}$.
III-B3 Tracking Controller
We formulate an MPC problem to track the optimized trajectory. Let $x=[p_{x},p_{y},\theta]^{T},u=[v,\omega]^{T}$, where $p_{x},p_{y}$ denote the $x,y$ coordinates of the vehicle, $\theta$ denotes the yaw angle, and $v,\omega$ denote the linear speed and angular rate. The optimization problem is then:
@@ III-B3 Tracking Controller
which is solved using do-mpc [13, 1, 45].
IV Simulation Experiments
@@
IV-A System Baselines
We evaluate TRAIL against two end-to-end navigation systems that include perception, planning, and control: Geometry and WayFASTER. These comparisons therefore reflect full navigation performance (success rate, execution time, etc.). For Geometry, we implement a method inspired by STEP [11] and RoadRunner [15]: geometric risk factors (step size, slope, curvature) are fused via Conditional Value at Risk (CVaR) to form a risk map, which is then used by an MPPI controller. Our implementation uses our encoder and the WayFASTER decoder [16] due to lack of publicly available code. We refer to this baseline as Geometry. For the traction-based baseline, we use WayFASTER [16] with its released network architecture [^2] https://github.com/matval/wayfaster, paired with an MPPI controller following the formulation in the original paper. We implemented both MPPI controllers by adapting a public PyTorch-based codebase [^3] https://github.com/UM-ARM-Lab/pytorch_mppi. We use a horzion $N=20$ with a time step of $dt=0.1$\mathrm{s}$\$.
Figure 3: Grassland scenario. The robot’s forward direction ($x$) and leftward direction ($y$) are represented by the red and green axes, respectively. Examples of paths produced by Geo-20 (blue), WF-20 (orange), and TRAIL (green) are shown on the left. The bottom left corner shows the camera image from the robot. On the right side, the top shows the initial path generated by A overlaid on the geometric risk map. Bottom right shows the optimized trajectory overlaid on the predicted bumpiness value. The trajectory is color-coded by speed.
IV-B Dataset and Training
@@ IV-C Metrics
IV-D Timing Analysis
Timings are performed on a workstation with an RTX 4090 GPU, an AMD Ryzen 9 7950x 16-core processor with 32 threads, and 128 GB of RAM. The encoder of our model and Geometry take $25$\mathrm{ms}$\$ per inference and WayFASTER’s encoder takes $20$\mathrm{ms}$\$. With horizon $N=20$, Geometry’s MPPI takes $10$\mathrm{ms}$\$ per inference and WayFASTER’s MPPI takes $25$\mathrm{ms}$\$ per inference. For the proposed method, A takes $5~$\mathrm{ms}$\$ (to generate a path with 40 points), trajectory optimization takes $7.5~$\mathrm{ms}$\$ per iteration (to optimize an initial path with 30 points, which are downsampled from the A path), and the MPC take $6$\mathrm{ms}$\$ with horizon $N=20$.
Figure 4: Forest scenario. Each row shows one set of experiments. The first column shows Path 1 and the second column shows Path 2 of each set of experiments. The forest consists of tall grass, bushes, rocks, logs, and trees. Images are tone- and color-corrected for visibility. The proposed method is able to navigate efficiently while the two baselines struggle: Geometry is conservative due to the density of objects with high elevations while WayFASTER is too optimistic since there are no collision events with similar objects in the training data. On the other hand, TRAIL is able to recognize and avoid rocks while not being constrained by the tall grass.
IV-E Navigation Performance
@@
Grassland The Grassland scenario (see Figure 3) features a rock hidden in tall grass (bottom left of Figure 3). The quantitative results in Table I show that Geometry, whose inflated geometric map cannot distinguish the rock from tall grass, produces an almost uniform risk map, drives straight toward the goal, runs over the rock, and incurs large vertical acceleration (Geo-20). Because CVaR values lie in $[0,1]$ [15], MPPI’s behavior is highly sensitive to the trade-off between risk and goal distance: high risk weight leads to oscillations and failure to reach the goal, while low risk weight reproduces Geometry’s straight-line behavior, consistent with our ablation study in Section IV-F. WayFASTER detects the rock but is overly optimistic, briefly clipping it and producing high vertical acceleration (WF-20), likely due to weak BEV cues from sparse supervision and a limited encoder. In contrast, our method maintains clearance, using bumpiness during trajectory optimization to push the path away from the rock while remaining less conservative in grass, yielding a smoother, shorter trajectory and an adaptive speed profile that slows near the rock and accelerates through grass. Increasing the MPPI horizon from $20$ to $50$ confirms that the baselines are not horizon-limited: Geo-50 becomes indecisive and slower due to accumulated risk, while WF-50 matches WF-20 (the single WF-20 failure resulted from odometry divergence). Thus, we use $N=20$ for subsequent experiments.
@@ IV-F Ablation Study
To further evaluate the proposed planner, we conduct an ablation study of different MPPI formualtions. We generate fixed-resolution cost maps from our trained model and using MPPI as the planner, following common practice. We implement five variants:
@@
In summary, each MPPI formulation minimizes the distance to goal and cost along the path, with a small penalty on control effort. MPPI-A-Bump has an additional cost term that penalizes deviations from the initial path. The planning horizon is set to $N=20$ with $dt=0.1s$ for all variants. Note that the smoothness terms in Equation 5 are not necessary for MPPI since dynamic feasibility is enforced by sampling bounded control inputs and rolling out the dynamics model.
@@
TABLE IV: Additional Simulation Investigation Results
Method Success (%) Progress (%) Time (s) Length (m) Acc (m/s2)
RMS Max
Grassland
MPPI-A*-Bump 100.0 100.0 $\pm$ 0.0 54.8 $\pm$ 4.6 15.5 $\pm$ 0.6 1.9 $\pm$ 0.2 41.6 $\pm$ 26.1
SALON 100.0 100.0 $\pm$ 0.0 27.8 $\pm$ 4.6 13.0 $\pm$ 2.6 2.1 $\pm$ 0.1 60.4 $\pm$ 42.3
SDF 100.0 100.0 $\pm$ 0.0 17.7 $\pm$ 1.4 13.9 $\pm$ 0.7 3.4 $\pm$ 1.1 53.4 $\pm$ 34.4
Forest Experiment I
Path 1 MPPI-A*-Bump 66.7 91.9 $\pm$ 11.5 19.7 $\pm$ 2.4 12.3 $\pm$ 1.3 2.3 $\pm$ 0.3 9.6 $\pm$ 4.1
SALON 0.0 64.4 $\pm$ 1.8 N/A N/A N/A N/A
SDF 100.0 100.0 $\pm$ 0.0 16.1 $\pm$ 4.1 14.6 $\pm$ 3.0 1.6 $\pm$ 0.1 5.3 $\pm$ 0.5
Path 2 MPPI-A*-Bump 66.7 76.6 $\pm$ 33.0 35.5 $\pm$ 17.2 14.9 $\pm$ 4.0 1.8 $\pm$ 0.4 5.6 $\pm$ 0.7
SALON 33.3 51.5 $\pm$ 34.7 17.00 $\pm$ 0.0 10.4 $\pm$ 0.0 1.9 $\pm$ 0.5 4.2 $\pm$ 0.5
SDF 66.7 66.7 $\pm$ 47.1 18.4 $\pm$ 0.3 17.0 $\pm$ 0.4 2.2 $\pm$ 0.4 5.4 $\pm$ 0.7
Forest Experiment II
Path 1 MPPI-A*-Bump 66.7 66.7 $\pm$ 47.1 36.5 $\pm$ 5.4 14.8 $\pm$ 1.7 3.2 $\pm$ 2.0 53.8 $\pm$ 66.0
SALON 0.0 20.0 $\pm$ 1.6 N/A N/A N/A N/A
SDF 66.7 98.0 $\pm$ 2.8 12.6 $\pm$ 0.7 10.7 $\pm$ 0.3 1.9 $\pm$ 0.1 6.8 $\pm$ 1.0
Path 2 MPPI-A*-Bump 100.0 100.0 $\pm$ 0.0 22.7 $\pm$ 4.2 13.0 $\pm$ 2.0 1.9 $\pm$ 0.2 5.5 $\pm$ 0.4
SALON 100.0 100.0 $\pm$ 0.0 28.2 $\pm$ 3.1 13.6 $\pm$ 0.8 1.7 $\pm$ 0.0 5.4 $\pm$ 0.5
SDF 66.7 97.6 $\pm$ 3.4 15.9 $\pm$ 0.9 14.0 $\pm$ 1.6 1.8 $\pm$ 0.2 5.4 $\pm$ 0.6
Forest Experiment III
Path 1 MPPI-A*-Bump 33.3 91.8 $\pm$ 5.8 16.2 $\pm$ 0.0 10.8 $\pm$ 0.0 2.0 $\pm$ 0.3 5.1 $\pm$ 0.6
SALON 0.0 55.6 $\pm$ 23.7 N/A N/A N/A N/A
SDF 33.3 89.9 $\pm$ 7.3 16.7 $\pm$ 0.0 14.4 $\pm$ 0.0 2.0 $\pm$ 0.1 6.2 $\pm$ 0.4
Path 2 MPPI-A*-Bump 33.3 58.6 $\pm$ 42.6 17.6 $\pm$ 0.0 11.7 $\pm$ 0.0 4.1 $\pm$ 2.4 129.2 $\pm$ 102.9
SALON 0.0 76.1 $\pm$ 3.4 N/A N/A N/A N/A
SDF 0.0 50.2 $\pm$ 31.4 N/A N/A N/A N/A
Forest Experiment IV
Path 1 MPPI-A*-Bump 33.3 61.6 $\pm$ 32.9 24.2 $\pm$ 0.0 12.6 $\pm$ 0.0 2.4 $\pm$ 0.7 54.4 $\pm$ 36.5
SALON 0.0 74.6 $\pm$ 2.3 N/A N/A N/A N/A
SDF 66.7 97.7 $\pm$ 3.2 14.8 $\pm$ 0.4 12.5 $\pm$ 0.5 2.2 $\pm$ 0.1 19.0 $\pm$ 2.9
Path 2 MPPI-A*-Bump 0.0 76.5 $\pm$ 11.9 N/A N/A N/A N/A
SALON 0.0 45.3 $\pm$ 36.6 N/A N/A N/A N/A
SDF 0.0 55.6 $\pm$ 10.6 N/A N/A N/A N/A
IV-G Additional Simulation Investigations
In addition to the end-to-end system baselines (Geometry, WayFASTER), we report simulation-only investigations to isolate specific factors: (i) the effect of global A initialization (MPPI-A-Bump), (ii) a trajectory-refinement variant that optimizes obstacle clearance/smoothness without bumpiness-driven speed adaptation (SDF with the same A initialization as TRAIL), and (iii) a small-data traversability learning comparison (SALON [41]), where we evaluate the released perception module paired with a custom MPPI controller due to the absence of a released planner formulation. The custom MPPI minimizes rollout cost from SALON’s cost map while enforcing the speed constraint derived from SALON’s speed map, consistent with the high-level description in [41]. Results are reported in table IV.
Overall, MPPI-A-Bump is sensitive to competing objective terms and exhibits oscillatory/indecisive behavior in our settings, leading to longer execution times (Table IV). SDF performs reasonably well in scenarios dominated by geometric obstacle avoidance, but degrades in cases involving tall vegetation, where obstacle geometry alone is ambiguous. SALON improves over WayFASTER in several small-data cases, but can miss hazards that are underrepresented in (or visually different from) its experience buffer, leading to failures in our forest scenarios. For example, SALON successfully assigns high cost to the rock region in Grassland, but can fail to detect leaf-covered rocks or trees in darker forest regions (e.g. Experiment IV Path 1 in Figure 4).
V Field Experiments
In hardware experiments, we focus on end-to-end system comparisons (Geometry and WayFASTER); simulation-only diagnostic studies (e.g., SALON, SDF, MPPI-A-Bump) are reported in Sec. IV-G. Our test platform is an AgileX Scout Mini equipped with a ZED 2i stereo camera and an Ouster OS1-32-GEN2.0 LiDAR. The training set contains fewer than 2000 samples collected in a similar area using a Clearpath Husky A300 UGV with the same sensor stack and calibration. The Husky data covers dirt, paved road, tall grass, and uneven terrain with roots and branches, without purposeful collisions with trees.
During data collection, Scout Mini odometry drifted substantially (despite post-processing), making label generation unreliable. The Husky’s smoother motion and traction yielded more stable odometry for offline aggregation and annotation. Training on Husky data and deploying on Scout Mini thus introduces a platform distribution shift (e.g., mass and vibration scale), providing an additional generalizability test. In practice, this shift did not noticeably degrade TRAIL: geometric outputs are expressed in the LiDAR frame, and thus transfer naturally across platforms when the sensor stack and calibration are unchanged, and planning relies mainly on relative variations in predicted bumpiness to shape the speed profile. We use KISS-ICP [44] for offline odometry in data processing and DLIO [9] for higher-rate odometry during deployment.